OpenStudy (anonymous):

Objects with masses m1 = 12.0 kg and m2 = 5.00 kg are connected by a light string that passes over a frictionless pulley. If, when the system starts from rest, m2 falls 1.00 m in 1.30 s, determine the coefficient of kinetic friction between m1 and the table.

OpenStudy (dape):

Have you drawn a free-body diagram?

OpenStudy (anonymous):

 work-kinetic energy \[W = DeltaEc\] m2*gh = 1/2m2*v^2 v^2 =2gh (1) use equation v = at (2) (1) and (2) \[\sqrt{2gh}=at\] a=3.4m/s^2 for m1: (ox) T - f =m1*a (3) for m2; (oy)m2*g -T = m2*a (4) take (3) + (4) m2*g - f =a (m1 + m2) f = m2*g - a (m1 + m2) f = coefficient of kinetic =f/m1*g